# Training Foundation Models as Data Compression: On Information, Model Weights and Copyright Law

### v1

@@ 1 Introduction
Besides curiosity and excitement, the current generative AI wave raises various philosophical and practical questions (Franceschelli & Musolesi, 2023; Shanahan, 2024; Weidinger et al., 2022). Among them, a relevant issue is how current copyright laws can be applied to generative AI (Lee et al., 2024). For instance, can AI-generated outputs be copyrighted? Is it lawful to train models on protected works? Furthermore, would it be possible to protect the model itself in some manner?
To better understand how such models work and how they can be interpreted under current laws, we believe that a viable solution is to link generative deep learning with information theory (Cover, 1999). In fact, generative models trained with self-supervised learning, i.e., by maximizing the likelihood of training data, as commonly done for foundation models (Bommasani et al., 2021) such as large language models (LLMs) (Bubeck et al., 2023; Gemini Team et al., 2023; Touvron et al., 2023) and diffusion models (Ho et al., 2020; Rombach et al., 2022), can be seen as a form of (lossy or lossless) compression (MacKay, 2003). From this perspective, the training algorithm plays the role of the compression algorithm; the inference (feed-forward) algorithm is the de-compression algorithm (with the input passed to the model working as a decoding key); and the model’s weights represent the compressed version of the training set.
@@
Deletang et al. (2024) discuss how a language model can implement a lossless compression process offline, i.e., through a fixed set of model parameters derived from training. We move a step further and claim the self-supervised learning used to train foundation models to be a lossy or lossless compression process, during which the whole training set is encrypted into the model’s weights. This is demonstrated by the fact that the model can reproduce certain portions of training samples (Carlini et al., 2023b; Kandpal et al., 2022). We suggest that the training makes the model’s weights the best possible compressed version of the training set or, more correctly, batches of it at a time. The analogies at the basis of the proposed training-as-compressing perspective are summarized in Figure 1. Building on this intuition, we envision the model’s weights as either a reproduction or a derivative work of training data. This new interpretation opens up a series of practical consequences that can be relevant from the copyright perspective. First, it provides a legal framework for understanding generative models which links them to the training data without impeding protection to the investment through the sui generis right. Then, it creates a direct link between the training data and the model outputs only made of steps requiring specific exceptions or authors’ authorization, thus allowing for potential requests for compensation.
@@ 2 Training-as-Compressing and Information
The propensity of foundation models to memorize and subsequently replicate training data is a topic that has received considerable attention in scholarly literature, as evidenced by works such as (Carlini et al., 2023a; Lee et al., 2022). In general, it is very hard to decompress every possible training sample perfectly and in its entirety, i.e., without any loss of information. Nonetheless, it has been shown that training samples can be retrieved (Carlini et al., 2021), and more advanced techniques might lead to an even higher degree of “retrievability”. The following experiment may help us understand this matter better.
In (1957), Noam Chomsky wrote the famous quote “Colorless green ideas sleep furiously” to demonstrate the distinction between syntax and semantics: the sentence is grammatically well-formed but semantically nonsensical. If a language model had learned the semantics of English, it should not generate a semantically nonsensical sentence, i.e., it should assign to semantically nonsensical words a small probability; if, on the contrary, such words are characterized by a large probability of being generated, then it is very likely that the model has memorized them. To test this, we use the same quote from Chomsky, and we check the probability of each subsequent word given the previous ones under an LLM (in our case, LLaMa3-70B (Meta, 2024)). The probability of ‘ green’ given ‘ “Colorless’ is 0.2, while for other colors like ‘ red’ or ‘ blue’ is 0.003. For all the subsequent words, i.e., ‘ ideas’, ‘ sleep’, and ‘ furiously’, the probability is always greater than 0.9. The model has essentially memorized the quote into its weights, otherwise it would have never assigned such high probabilities to a semantically nonsensical sentence.
@@
From an information theoretic perspective, such training data compression can be explained through the information bottleneck (IB) principle (Tishby et al., 1999). The IB principle applies when we aim to extract relevant information from an input variable $X\in\mathcal{X}$ about an output variable $Y\in\mathcal{Y}$. Given their joint distribution $p(X,Y)$, the relevant information is defined as the mutual information $I(X;Y)$. With $\hat{X}$ as the relevant part of $X$ with respect to $Y$, the IB method aims to find the optimal $\hat{X}\in\mathcal{\hat{X}}$, i.e., the one that minimizes $I(X;\hat{X})$ (obtaining the simplest possible statistics) while maximizing $\beta I(\hat{X};Y)$ (containing all the relevant information). Tishby and Zaslavsky (2015) argued that neural networks could be interpreted under the theoretical framework of the IB principle. Indeed, neural networks learn to extract efficient representations of the relevant features $\hat{X}$ of the input $X$ for predicting the output $Y$, given a finite sample of the joint distribution $p(X,Y)$. In the context of supervised learning, this means ignoring the irrelevant part of $X$ by only selecting the one needed to predict $Y$. However, in self-supervised learning $Y\approx X$. It follows that $\hat{X}$ is the relevant part of $X$ with respect to itself, so it is a compressed version of $X$.
@@ 4 Implications
First, it provides a legal framework to understand them, removing the veil of uncertainty surrounding this issue. Although asserting copyright protection for weights as a derivative work presents challenges due to the absence of valid authorship (Otero, 2021), it is possible to safeguard them by viewing the file with the model’s weights as a database. Indeed, they can be considered as a collection of floating point numbers, which can be retrieved independently. Moreover, the significant investments required for obtaining them make the model’s weights eligible for the sui generis right (thus providing certain rights to those who have invested in the database constitution independently from its copyright protection) (Sousa e Silva, 2024). In other words, the sui generis right can protect the investment; our copyright perspective can link the model’s weights back to the training data, providing a new perspective over one of the several issues concerning the generative-AI supply chain (Lee et al., 2024). The same considerations still hold in the case of a fine-tuned model. According to Lee et al. (2024), this would be considered a derivative work of the pre-trained model (and also of the fine-tuning data). In other words, fine-tuning could be considered as nothing more than an additional step in the information processing chain. Again, the weights of the fine-tuned model would be eligible for the sui generis right. However, whether it qualifies for protection as a derivative work remains an open question, and the determination of valid (human) authorship can vary on a case-by-case basis.
@@
Second, this type of interpretation provides a potential framework for works generated by the model. Indeed, decompressing the information from the model might be seen as producing a derivative work of the weights, thus a derivative work of a copy of a protected work or a derivative work of a derivative work of (a copy of) a protected work. Either way, this link between the output and the training data may help enforce their copyrights. It is worth noting that the EU text and data mining (TDM) exceptions (European Union, 2019) as well as other comparable rules (Fiil-Flynn et al., 2022) apply for TDM purposes, such as training the model, therefore to the case of the creation of a copy or derivative work; however, they do not apply for further derivative works from the model. A similar consideration can also be drawn for the US fair use doctrine, which arguably applies to training a model on copyrighted data but is less likely when deployed to generate similar content that can threaten their market (Henderson et al., 2023). The main consequence is that authorization from the training set’s rightsholders would be required (or else the reproduction or adaptation right would be triggered), allowing for potential requests for compensation from original authors. In addition, generated works would need to respect the moral rights of the owners of training data, even when their economic rights have expired. The fact that a new derivative work is protected by copyright is an entirely different issue that will depend on the human contribution (i.e., the input to the model), in particular, on its substantiality and its being the main contribution of the originality (Guadamuz, 2017; Franceschelli & Musolesi, 2022). The overall conceptual framework based on the proposed training-as-compressing perspective is summarized in Figure 2.
@@ 5.1 Deep Learning and Information Theory
Information theory is underpinning machine learning (MacKay, 2003). Recently, there has been considerable interest in the intersection between compression and neural networks, either to learn efficient data compression algorithms (Goyal et al., 2019; Yang et al., 2023) or to compress the learned models themselves (Choudhary et al., 2020; Polino et al., 2018). The problem of training data compression has been analyzed from the point of view of the information bottleneck principle (IB) (Tishby et al., 1999), which provides a methodology for extracting a quantitative measurement of “meaningful” and “relevant” information. In particular, IB can be used to derive an optimal theoretical limit for training neural networks in terms of trade-off between compression and prediction (Tishby & Zaslavsky, 2015). From this perspective, stochastic gradient descent can be seen as composed of two distinct phases: a “fitting” phase in which the mutual information between the weights and the target output increases and a “compression” phase in which the mutual information between the weights and the input decreases (Shwartz-Ziv & Tishby, 2017). However, the latter phenomenon might not be an intrinsic feature of deep neural networks, but only the result of the presence of nonlinearities in the networks (Saxe et al., 2018). An important recent paper in this context is (2024), in which Schwartz-Ziv and LeCun discuss IB in relation to self-supervised learning, especially in the context of multimodal learning; this work is particularly relevant given the prominence of these techniques in training state-of-the-art foundation models.
@@ 5.2 Generative AI and Copyright
There has been a long-standing interest in the copyright issues around generative AI (Butler, 1982; Samuelson, 1986). Different legal issues are at play when considering the entire generative-AI supply chain (Lee et al., 2024). Whether training neural networks on protected data is lawful has been highly debated across different national legislations (Lemley & Casey, 2021), e.g., U.S. fair use (Henderson et al., 2023; Sobel, 2017), EU text and data mining exceptions (Geiger et al., 2018; Sobel, 2021), and others (Guadamuz, 2024; Samuelson, 2021). The other main debate has focused on whether a machine-generated work is protected by copyright or not (Craig & Kerr, 2019; Gervais, 2020; Grimmelman, 2016) and on the question of who might own its ownership in the future (Bonadio & McDonagh, 2020; Franceschelli & Musolesi, 2022). However, other issues have also been considered, such as whether the model output can infringe the reproduction right (Gillotte, 2020; Vyas et al., 2023) or how the trained model can be protected (Otero, 2021; Slowinski, 2021). Our analysis mainly concerns the latter point, but also provides a new technical understanding and perspective on the copyright of AI-generated works.

### v2

@@ 1 Introduction
Besides curiosity and excitement, the current generative AI wave raises various philosophical and practical questions (Shanahan, 2024; Weidinger et al., 2022). Among them, a relevant issue is how current copyright laws can be applied to generative AI (Lee et al., 2024). For instance, can AI-generated outputs be copyrighted? Is it lawful to train models on protected works? Furthermore, would it be possible to protect the model itself in some manner?
To better understand how such models work and how they can be interpreted under current laws, we believe that a viable solution is to link generative deep learning with information theory (Cover, 1999). In fact, generative models trained with self-supervised learning, i.e., by maximizing the likelihood of training data, as commonly done for foundation models (Bommasani et al., 2021), such as large language models (LLMs) (Gemini Team et al., 2023; Meta, 2024) and diffusion models (Rombach et al., 2022), can be seen as a form of (lossy or lossless) compression (MacKay, 2003). From this perspective, the training algorithm plays the role of the compression algorithm; the inference (feed-forward) algorithm is the de-compression algorithm (with the input passed to the model working as a decoding key); and the model’s weights represent the compressed version of the training set.
@@
Deletang et al. (Deletang et al., 2024) discuss how a language model can implement a lossless compression process offline, i.e., through a fixed set of model parameters derived from training. We move a step further and claim the self-supervised learning used to train foundation models to be a lossy or lossless compression process, during which the whole training set is encrypted into the model’s weights. This is demonstrated by the fact that the model can reproduce certain portions of training samples (Carlini et al., 2023b). We suggest that the training optimizes the model’s weights to be the best possible compressed version of the training set, or more accurately, batches of it at a time. The analogies at the basis of the proposed training-as-compressing perspective are summarized in Figure 1. Building on this intuition, we envision the model’s weights as either a reproduction or a derivative work of training data. This new interpretation opens up a series of practical consequences that can be relevant from the copyright perspective. First, it provides a legal framework for understanding generative models which links them to the training data without impeding protection to the investment through the sui generis right. Then, it creates a direct link between the training data and the model outputs only made of steps requiring specific exceptions or authors’ authorization, thus allowing for potential requests for compensation.
@@ 2 Training-as-Compressing and Information
The propensity of foundation models to memorize and subsequently replicate training data is a topic that has received considerable attention in scholarly literature, as evidenced by works such as (Carlini et al., 2023a). In general, it is very hard to decompress every possible training sample perfectly and in its entirety, i.e., without any loss of information. Nonetheless, it has been shown that training samples can be retrieved (Carlini et al., 2021), and more advanced techniques might lead to an even higher degree of “retrievability”. The following experiment may help us understand this matter better.
In 1957, Noam Chomsky introduced the famous sentence “Colorless green ideas sleep furiously” to demonstrate the distinction between syntax and semantics: the sentence is grammatically well-formed but semantically nonsensical. If a language model had learned the semantics of English, it should not generate a semantically nonsensical sentence, i.e., it should assign to semantically nonsensical words a small probability; if, on the contrary, such words are characterized by a large probability of being generated, then it is very likely that the model has memorized them. To test this, we use the same quote from Chomsky, and we check the probability of each subsequent word given the previous ones under an LLM (in our case, LLaMa3-70B (Meta, 2024)). As depicted in Figure 2, the probability of ‘ green’ given ‘ “Colorless’ is 0.2, while for other colors like ‘ red’ or ‘ blue’ is 0.003. For all the subsequent words, i.e., ‘ ideas’, ‘ sleep’, and ‘ furiously’, the probability is always greater than 0.9. The model has essentially memorized the quote into its weights, otherwise it would have never assigned such high probabilities to a semantically nonsensical sentence.
Figure 2: A thought experiment to confirm that LLMs memorize training data: even if the sentence is semantically nonsensical, the model assigns high probability to its tokens just because it occurred in its training set.
@@
From an information theoretic perspective, such training data compression can be explained through the information bottleneck (IB) principle (Tishby et al., 1999). The IB principle applies when we aim to extract relevant information from an input variable $X\in\mathcal{X}$ about an output variable $Y\in\mathcal{Y}$. Given their joint distribution $p(X,Y)$, the relevant information is defined as the mutual information $I(X;Y)$. With $\hat{X}$ as the relevant part of $X$ with respect to $Y$, the IB method aims to find the optimal $\hat{X}\in\mathcal{\hat{X}}$, i.e., the one that minimizes $I(X;\hat{X})$ (obtaining the simplest possible statistics) while maximizing $\beta I(\hat{X};Y)$ (containing all the relevant information). Tishby and Zaslavsky (Tishby & Zaslavsky, 2015) argued that neural networks could be interpreted under the theoretical framework of the IB principle. Indeed, neural networks learn to extract efficient representations of the relevant features $\hat{X}$ of the input $X$ for predicting the output $Y$, given a finite sample of the joint distribution $p(X,Y)$. In the context of supervised learning, this means ignoring the irrelevant part of $X$ by only selecting the one needed to predict $Y$. However, in self-supervised learning $Y\approx X$. It follows that $\hat{X}$ is the relevant part of $X$ with respect to itself, so it is a compressed version of $X$.
@@ 4 Implications
First, it provides a legal framework to understand them, removing the veil of uncertainty surrounding this issue. Although asserting copyright protection for weights as a derivative work presents challenges due to the absence of valid authorship (Otero, 2021), it is possible to safeguard them by viewing the file with the model’s weights as a database. Indeed, they can be considered as a collection of floating point numbers, which can be retrieved independently. Moreover, the significant investments required for obtaining them make the model’s weights eligible for the sui generis right (thus providing certain rights to those who have invested in the database constitution independently from its copyright protection) (Sousa e Silva, 2024). In other words, the sui generis right can protect the investment; our copyright perspective can link the model’s weights back to the training data, providing a new perspective over one of the several issues concerning the generative-AI supply chain (Lee et al., 2024). The same considerations still hold in the case of a fine-tuned model. According to Lee et al. (Lee et al., 2024), this would be considered a derivative work of the pre-trained model (and also of the fine-tuning data). In other words, fine-tuning could be considered as nothing more than an additional step in the information processing chain. Again, the weights of the fine-tuned model would be eligible for the sui generis right. However, whether it qualifies for protection as a derivative work remains an open question, and the determination of valid (human) authorship can vary on a case-by-case basis.
@@
Second, this type of interpretation provides a potential framework for works generated by the model. Indeed, decompressing the information from the model might be seen as producing a derivative work of the weights, thus a derivative work of a copy of a protected work or a derivative work of a derivative work of (a copy of) a protected work. Either way, this link between the output and the training data may help enforce their copyrights. It is worth noting that the EU text and data mining (TDM) exceptions as well as other comparable rules (Fiil-Flynn et al., 2022) apply for TDM purposes, such as training the model, therefore to the case of the creation of a copy or derivative work; however, they do not apply for further derivative works from the model. A similar consideration can also be drawn for the US fair use doctrine, which arguably applies to training a model on copyrighted data but is less likely when deployed to generate similar content that can threaten their market (Henderson et al., 2023). The main consequence is that authorization from the training set’s rightsholders would be required (or else the reproduction or adaptation right would be triggered), allowing for potential requests for compensation from original authors. In addition, generated works would need to respect the moral rights of the owners of training data, even when their economic rights have expired. The fact that a new derivative work is protected by copyright is an entirely different issue that will depend on the human contribution (i.e., the input to the model), in particular, on its substantiality and its being the main contribution of the originality (Guadamuz, 2017; Franceschelli & Musolesi, 2022). Crucially, these considerations also apply to synthetic data (i.e., AI-generated works) used as new training data for a different foundation model (e.g., (Shumailov et al., 2024)). Essentially, the chain of relationships between the resulting model-generated data and the original protected training data would be longer, but it would still involve steps that trigger either the reproduction or adaptation rights. The overall conceptual framework based on the proposed training-as-compressing perspective is summarized in Figure 3.
@@ 5.1 Deep Learning and Information Theory
Information theory is underpinning machine learning (MacKay, 2003). Recently, there has been considerable interest in the intersection between compression and neural networks, either to learn efficient data compression algorithms (Yang et al., 2023) or to compress the learned models themselves (Choudhary et al., 2020). The problem of training data compression has been analyzed from the point of view of the information bottleneck principle (IB) (Tishby et al., 1999), which provides a methodology for extracting a quantitative measurement of “meaningful” and “relevant” information. In particular, IB can be used to derive an optimal theoretical limit for training neural networks in terms of trade-off between compression and prediction (Tishby & Zaslavsky, 2015). From this perspective, stochastic gradient descent can be seen as composed of two distinct phases: a “fitting” phase in which the mutual information between the weights and the target output increases and a “compression” phase in which the mutual information between the weights and the input decreases (Shwartz-Ziv & Tishby, 2017). However, the latter phenomenon might not be an intrinsic feature of deep neural networks, but only the result of the presence of nonlinearities in the networks (Saxe et al., 2018). An important recent paper in this context is (Shwartz-Ziv & LeCun, 2024), in which Schwartz-Ziv and LeCun discuss IB in relation to self-supervised learning, especially in the context of multimodal learning; this work is particularly relevant given the prominence of these techniques in training state-of-the-art foundation models.
@@ 5.2 Generative AI and Copyright
There has been a long-standing interest in the copyright issues around generative AI (Butler, 1982; Samuelson, 1986). Different legal issues are at play when considering the entire generative-AI supply chain (Lee et al., 2024). Whether training neural networks on protected data is lawful has been highly debated across different national legislations (Lemley & Casey, 2021), e.g., U.S. fair use (Henderson et al., 2023), EU text and data mining exceptions (Sobel, 2021), and others (Guadamuz, 2024; Samuelson, 2021). The other main debate has focused on whether a machine-generated work is protected by copyright (Craig & Kerr, 2019; Grimmelman, 2016), and who might own its copyright in the future (Franceschelli & Musolesi, 2022). However, other issues have also been considered, such as whether the model output can infringe the reproduction right (Vyas et al., 2023) or how the trained model can be protected (Otero, 2021). Our analysis mainly concerns the latter point, but also provides a new technical understanding and perspective on the copyright of AI-generated works.